During one grading period your test scores are 83, 79, 92, 86, 93
What is the mean, mode, and median
Which measure of central tendency would give you the highest grade

Answer by venugopalramana(3286)

ARRANGE THE SCORES IN INCREASING ORDER...WE GET
79,83,86,92,93
NOW
MEAN =(79+83+86+92+93)/5=433/5=86.6
MEDIAN=MIDLE SCORE=86
MODE=SCORE WITH MAXIMUM FREQUENCY...HERE ALL HAVE SAME FREQUENCY ...ONE ONLY ...IN CASE 2 OR MORE SCORES HAVE SAME FREQUENCY WE TAKE THEIR AVERAGE AS THE MODE...HERE MODE COINCIDES WITH MEAN AS ALL SCORES ARE HAVING SAME FREQUENCY.
MEAN IS CONSIDERED THE BEST REPRESENTATION FOR CENTRAL TENDENCY.